HanAkoSolo (jamierox4ev3r):

Quick question (I just want someone to check my answer) If the standard form of a parabola with a vertex (h,k) is (x-h)^2 = 4p(y-k) then simplify this equation into standard form: x^2 +2x-y+3=0

HanAkoSolo (jamierox4ev3r):

If I'm correct, this is a problem on completing the square

HanAkoSolo (jamierox4ev3r):

so this is what I've done so far: \[x ^{2}+2x-y+3=0\] \(x^{2}+2x+1=y-3+1\) this then becomes: \((x-1)^{2}\) = \(1(y-2)\)

HanAkoSolo (jamierox4ev3r):

Is this literally all i need to do? o3o

HanAkoSolo (jamierox4ev3r):

@sleepyjess @e.mccormick

OpenStudy (anonymous):

a bit correction.. it should be x+1 and to look it more standard write it as (x+1)^2 = 4*1/4(y-2)

HanAkoSolo (jamierox4ev3r):

oops yeah i had addition instead of subtraction, that was just a typo. My mistake.

HanAkoSolo (jamierox4ev3r):

And oh i see why you would make those additions, since p is muliplied by 4...are they really necessary additions though? @divu.mkr ?

OpenStudy (anonymous):

yep